Circular No. 7061

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ew IN NGC 6943 J. Maza, Department of Astronomy, University of Chile, reports the

discovery by R. Antezana, on a T-Max 400 film taken by L. E. González with the Maksutov telescope at Cerro El Roble on Nov. 23.02 UT, of a supernova (B ∼ 14) located at α = 20^{h}44^{m}32^{s}.1, δ = −68^{o}43′37′′ (equinox 2000.0), which is 5′′ west and 73′′ north of the nucleus of NGC 6943. The object was confirmed by R. González on Nov. 24.07 with the same telescope at Cerro El Roble. SN 1998ew is not visible on films obtained with the Cerro El Roble telescope on Nov. 11.06 (B > 19.5).

F. Patat, European Southern Observatory (ESO); N. Christlieb, Hamburger Sternwarte; and J. Sollerman, Stockholm Observatory, report that a full reduction of a CCD spectrogram (range 400–900 nm, resolution 2.0 nm) obtained with the ESO-Danish 1.54-m telescope (+ DFOSC) on Nov. 27.05 UT shows that a few lines, accompanied by very weak P-Cyg profiles, are superimposed on a blue and otherwise featureless continuum, and they are identified as Hα, Hβ, Hγ, and He I at 587.6 nm. The spectral features and the blue continuum are consistent with a supernova of type II, roughly around maximum light. The expansion velocities deduced from the minima of Hβ, Hγ, and He 587.6-nm are 8200, 8500, and 7900 km/s, respectively. The magnitudes of SN 1998ew were V ∼ 15.2 and R ∼ 14.9. Using a recession velocity of 3085 km/s for NGC 6943 (Da Costa et al. 1991, Ap.J. Suppl. 75, 935) and H_{0} = 65 km s^{−1} Mpc^{−1}, the absolute magnitude is V ∼ −18.2, which is somewhat brighter than that seen in normal type-II supernovae. There may be some resemblance to bright type-II supernovae, such as SN 1979C (Branch et al. 1981, Ap.J. 224, 780).

SUPERNOVA 1998ex IN MCG +11-10-16 J. Mueller reports her discovery of a supernova (mag ∼ 18.5) on a red

plate taken by herself with the 1.2-m Oschin Schmidt Telescope on Nov. 25 UT in the course of the second Palomar Sky Survey. SN 1998ex is located at α = 7^{h}41^{m}42^{s}.28, δ = +64^{o}44′02′′.3 (equinox 2000.0), which is 5′′.1 west and 10′′.1 north of the center of MCG +11-10-16. R. Gal obtained CCD images of SN 1998ex with the 1.5-m Palomar telescope on Nov. 26, and he provided the positional information. There is no object at this position on either the Digital Sky Survey or on an overlapping red sky-survey plate taken on 1989 Dec. 23. A spectrum taken on Nov. 26 by I. Smail and A. Dressler with the Hale 5-m telescope shows it to be a type-Ia supernova. A preliminary measurement of the redshift of the supernova is 11 200 km/s.

1998 November 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green